# 3D variational autoencoder for fingerprinting microstructure volume elements

### v2

@@ Abstract
Microstructure quantification is an important step towards establishing structure-property relationships in materials. Machine learning-based image processing methods have been shown to outperform conventional image processing techniques and are increasingly applied to microstructure quantification tasks. In this work, we present a 3D variational autoencoder (VAE) for encoding microstructure volume elements (VEs) comprising voxelated crystallographic orientation data. Crystal symmetries in the orientation space are accounted for by mapping to the crystallographic fundamental zone as a preprocessing step, which allows for a continuous loss function to be used and improves the training convergence rate. The VAE is then used to encode a training set of VEs with an equiaxed polycrystalline microstructure with random texture. Accurate reconstructions are achieved with a relative average misorientation error of $9\times 10^{-3}$, on the test dataset, for a continuous latent space with dimension 256. We show that the model generalises well to microstructures with textures, grain sizes and aspect ratios outside the training distribution.
Structure-property relationships are explored through using the training set of VEs as initial configurations in various crystal plasticity (CP) simulations. Microstructural fingerprints extracted from the VAE, which parameterise the VEs in a low-dimensional latent space, are stored alongside the volume-averaged stress response, at each strain increment, to uniaxial tensile deformation from CP simulations. This is then used to train a fully connected neural network mapping the input fingerprint to the resulting stress response, which acts as a surrogate model for the CP simulation. The fingerprint-based surrogate model is shown to accurately predict the microstructural dependence in the CP stress response, with a relative mean-squared error (MSE) of $8.9\times 10^{-4}$ on unseen test data. This approach offers a significant speed-up on the order of $10^{6}$ for a stress-strain curve prediction, compared to running a CP simulation.
@@ 1 Introduction
To understand the macroscopic behaviour of materials beyond the regimes that have been probed experimentally, it is important to consider structure-property relations from the atomistic scale up through the microstructural and continuum length scales [1, 2, 3]. Addressing this key challenge is hindered by the severe computational costs associated with conventional concurrent [4] and coarse-grained multiscale approaches [5]. More recently, deep learning-based methods have shown great promise in emulating coarse-grained reduced order models whilst offering significant computational speed-up [6, 7]. At the microstructural scale, deep learning-based reduced order models for crystal plasticity (CP) have been proposed, which utilise representative volume elements for a given material, and predict homogenized stress responses for given load paths [8, 9, 10, 11]. While these approaches have been shown to provide accurate predictions, training is often limited to smaller volume elements (VEs), that are not representative of the microstructure distribution. As a result, such model responses can have a much higher variance and becomes heavily microstructure dependent.
@@
To address this task, a variety of methods have been proposed over the last few decades. Local descriptors, centred at key points in an image, can be aggregated with statistical methods to construct distributions of typical features that can quantify a microstructure [23, 24, 25] and this has been shown to provide a useful parameterisation for classification tasks. Linear projection, through principal component analysis, has also been used to develop physically meaningful descriptors [26]. More recently, attention has shifted towards convolutional neural networks (CNNs) [27]. Pretrained networks for popular architectures, such as AlexNet [28], are readily available and have been shown to adapt well to microstructure, despite having been trained on natural images [29]. Fine-tuning can then be applied to further optimise a pretrained network for a specific task, such as constructing structure-property relationships [30]. One of the drawbacks with a standard CNN is the requirement for a predefined task and labelled data to train the network. On the other hand, an unsupervised method for training a CNN is the variational autoencoder (VAE) [31]. VAEs are composed of a pair of CNNs, one for encoding the input and another that aims to reconstruct the input from the encoded representation. Once trained, the encoder portion of the VAE has been shown to provide suitable microstructure parameterisation for structure-property predictions [32, 33].
@@
The trained VAE is then used to generate a dataset of microstructural fingerprints, and DAMASK [39] is used to perform CP simulations of uniaxial tension applied to the corresponding VEs. The resulting fingerprints and volume-averaged stress responses at each strain increment are then used to train and validate a fully connected neural network surrogate for the CP model. There are a wide range of options for the surrogate model, but a fully connected architecture was selected as there are no limitations on suitable input size (unlike Gaussian process regression). The time-dependence of the loading is implicit in the model, as this is constant for each simulation, so a fixed number of output nodes could be used. To extend this work to operate on an arbitrary load path, recurrent neural networks (RNNs) and/or transformers would need to be considered. Therefore, the full pipeline consists of VE generation, VAE fingerprinting, CP simulations of uniaxial tension and, finally, a fully connected network to predict stress response. The latent space constructed by the trained VAE is explored with dimensionality reduction and spectral clustering to interpret how various metrics are distributed across the space. We show that fingerprinting microstructure VEs in this way can provide a suitable descriptor for parameterising microstructure as inputs for a microstructure-dependent CP surrogate model.
@@ 2.1 Datasets
Several datasets of 3D single phase VEs were constructed using Dream3D [40]. Each VE is converted into a PyTorch [41] tensor with shape $(3,64,64,64)$, i.e. a $64^{3}$ VE with 3 channels at each voxel to describe crystal orientation as an Euler angles in the Bunge convention. The Euler angles are normalised to a range of $[0,1]$ for training purposes, by dividing by $(2\pi,\pi,2\pi)$ across the three channels respectively. We also use the DAMASK Python library [42] to map the Euler angles into the crystallographic fundamental zone to remove crystal symmetries as a preprocessing step. This requires prior knowledge about the crystal structure. We consider cubic and hexagonal symmetries here and, as such, the trained models do not extend beyond that. Most of the results presented herein use a cubic crystal structure. The hexagonal structure is introduced for the surrogate model, as it provides a more anisotropic stress response to train against. There are several options for taking symmetry into account, but it was deemed that addressing crystallographic symmetries during preprocessing provided the most stability during training. Alternatively, symmetries could be considered in the loss function during training. However, we found that this caused numerical instabilities, due to enforcing a discontinuous loss function, and it can be expensive to compute at each epoch.
2.1.1 Training Dataset
@@ 2.2.1 Variational Autoencoders (VAEs)
$z_{i}=\mu_{i}+e^{\frac{1}{2}\ln(\sigma_{i})\mathcal{N}(0,1)},$
@@ 2.2.3 Loss Functions
$\mathcal{L}_{\text{FFT}}(x,\hat{x})=\frac{1}{m}\sum_{j=1}^{m}\left[(\text{Im}\{\mathcal{F}(x)_{j}\}-\text{Im}\{\mathcal{F}(\hat{x})_{j}\})^{2}+(\text{Re}\{\mathcal{F}(x)_{j}\}-\text{Re}\{\mathcal{F}(\hat{x})_{j}\})^{2}\right],$
@@ 2.3 Crystal Plasticity (CP) Model
The open source CP simulation software DAMASK [39] was used alongside MatFlow [46] for managing high-concurrency workflows to generate the training data for the surrogate model. The workflow generates a VE using Dream3D, as described in Section 2.1.1, post-processes the output into a format readable by DAMASK and performs a CP simulation with uniaxial loading along the $x$-axis. Each VE is deformed to a strain of 0.01 at a rate of $1\times 10^{-3}$ over 20 increments, to characterize the plastic yielding behaviour of the VE. The VE is also passed through the trained VAE to construct a microstructural fingerprint in parallel with performing the CP simulation. Finally, the workflow output is processed to yield a mapping between each VE’s fingerprint, and the associated volume-averaged stress at each strain increment. The workflow was executed to run 50,000 simulations, 48,401 of which converged to a solution, with some runs not converging due to the severe plastic anisotropy of HCP magnesium. This was split into a training set of 40,000, with the rest reserved for model validation. Each simulation was run across 4 cores on Intel Sapphire Rapids CPUs, with a wall-clock time of $\sim 10$ minutes per simulation. The volume-averaged stress responses were normalised using min-max scaling, across the whole dataset, to a range of [0, 1], prior to splitting the dataset into training and validation partitions. Single crystal parameters for a magnesium phenomenological CP model [47] were used to generate this dataset, as it provides a relatively large spread in stress response compared to copper due to higher anisotropy. The anisotropy comes from magnesium having a HCP crystal structure, and so the VAE was retrained, in the same manner as previously described and with the relevant fundamental zone, before being used for fingerprint construction.
@@ 2.4 Surrogate Model
A fully connected neural network was constructed as a surrogate for the CP simulations discussed in Section 2.3. The network has an input layer size that matches the dimensionality of the fingerprints being operated on. This is followed by a hidden layer of size 512, with Leaky ReLU activation, and a final 21-dimensional layer, with sigmoid activation, to output the normalised stress response for each discrete strain increment. The Adam optimiser is used here again, operating on the mean-squared error between the prediction and true stress from the DAMASK simulations. A learning rate of $1\times 10^{-5}$ was used with a weight decay of $1\times 10^{-5}$ and a batch size of 128 was used to introduce some stochasticity into the training process, rather than backpropagating after a forward pass with the entire training set (which would have fit into memory with the resources available). This stochasticity is favourable here, as the training dataset is relatively small.
@@ 3.1.2 Network Depth Selection
ResNet12 $0.153\pm 0.025$ $0.153\pm 0.025$
ResNet18 $0.008\pm 0.005$ $0.009\pm 0.006$
ResNet34 $0.052\pm 0.016$ $0.052\pm 0.016$
@@
Despite the drop in BCE loss from ResNet18 to ResNet34, the mean misorientation is an order of magnitude lower for ResNet18. This confirms that, although the spectral term in the combined loss function is proposed as a method for regularising the latent space [45], it does also act as a reconstruction loss as well. Further analysis of the latent space, to confirm that is has been suitably regularised, is presented in Section 3.3.
@@ 3.2 Reconstruction Accuracy
The large and elongated grains are reconstructed with an accuracy comparable to the training data, as shown in Table 5. However, reconstruction in the case of the small grains is clearly not representative of the input VE. This could be a caveat of the minimum resolution of a VAE as, due to the inherently smooth outputs from the convolution transpose operations, there is a minimum resolution that can be achieved. The grain size distributions used to construct the datasets used here could be interpolated to hone in on the value of this minimum threshold, but that is beyond the scope of this work. Alternatively, this could be due to a lack in representation of small grains in the training data and issues arising with extrapolating to accommodate these. Extrapolation up to the large grains and elongated grains from the training data is likely feasible due to these morphologies being obtainable by placing grains from the training data with equal orientations adjacent to each other.
Morphology Mean misorientation
Large grains $0.006\pm 0.007$
Small grains $0.072\pm 0.107$
Elongated grains $0.009\pm 0.009$
Table 5: Mean misorientations between input VEs and their corresponding reconstructions for 100 samples from each of the morphological feature edge cases.
@@
As before, the mean misorientation was calculated across 100 samples for each of these edge cases and the results are provided in Table 6. The reconstruction accuracy here is comparable to the training data for the brass and copper examples. However, there is a significant improvement over the training loss for the cube and Goss textures.
@@
Brass $6\times 10^{-3}\pm 3\times 10^{-3}$
Copper $6\times 10^{-3}\pm 4\times 10^{-4}$
Cube $9\times 10^{-4}\pm 3\times 10^{-5}$
Goss $7\times 10^{-4}\pm 3\times 10^{-5}$
Table 6: Mean misorientations between input VEs and their corresponding reconstructions for 100 samples from each of the textured edge cases.

### v3

@@ Abstract
Microstructure quantification is an important step towards establishing structure-property relationships in materials. Machine learning-based image processing methods have been shown to outperform conventional image processing techniques and are increasingly applied to microstructure quantification tasks. In this work, we present a 3D variational autoencoder (VAE) for encoding microstructure volume elements (VEs) comprising voxelated crystallographic orientation data. Crystal symmetries in the orientation space are accounted for by mapping to the crystallographic fundamental zone as a preprocessing step, which allows for a continuous loss function to be used and improves the training convergence rate. The VAE is then used to encode a training set of VEs with an equiaxed polycrystalline microstructure with random texture. Accurate reconstructions are achieved with a relative average misorientation error of $3\times 10^{-2}$, on the test dataset, for a continuous latent space with dimension 256. We show that the model generalises well to microstructures with textures, grain sizes and aspect ratios outside the training distribution.
The main aim of this paper is to present the 3D VAE as an application-agnostic method for parameterising microstructure for input into downstream tasks where microstructure dependence is required. As a proof of concept, a simple surrogate model for uniaxial crystal plasticity (CP) simulations, with a fixed load path and microstructural dependence is presented. Microstructural fingerprints, obtained by encoding VEs with the trained VAE encoder, parameterise the VEs in a low-dimensional latent space and are stored alongside the volume-averaged stress response, at each strain increment. This is then used to train a fully connected neural network mapping the input fingerprint to the resulting stress response, which acts as a surrogate model for the CP simulation. The fingerprint-based surrogate model is shown to accurately predict the microstructural dependence in the CP stress response, with a mean relative error (MSE) of 2.75 MPa on unseen test data. This approach offers a significant speed-up on the order of $10^{8}$ for a stress-strain curve prediction, compared to running a CP simulation.
@@ 1 Introduction
To understand the macroscopic behaviour of materials beyond the regimes that have been probed experimentally, it is important to consider structure-property relations from the atomistic scale up through the microstructural and continuum length scales [1, 2, 3]. Addressing this key challenge is hindered by the severe computational costs associated with conventional concurrent [4] and coarse-grained multiscale approaches [5]. More recently, deep learning-based methods have shown great promise in emulating coarse-grained reduced order models whilst offering significant computational speed-up [6, 7]. At the microstructural scale, deep learning-based reduced order models for crystal plasticity (CP) have been proposed, which utilise representative volume elements for a given material, and predict homogenized stress responses for given load paths [8, 9, 10, 11]. While these approaches have been shown to provide accurate predictions, training is often limited to smaller volume elements (VEs), that are not representative of the microstructure distribution [12]. As a result, such model responses can have a much higher variance and become heavily microstructure dependent.
@@
To address this task, a variety of methods have been proposed over the last few decades. Local descriptors, centred at key points in an image, can be aggregated with statistical methods to construct distributions of typical features that can quantify a microstructure [24, 25, 26] and this has been shown to provide a useful parameterisation for classification tasks. Linear projection, through principal component analysis (PCA), has also been used to develop physically meaningful descriptors [27]. More recently, attention has shifted towards convolutional neural networks (CNNs) [28]. Pretrained networks for popular architectures, such as AlexNet [29], are readily available and have been shown to adapt well to microstructure, despite having been trained on natural images [30]. Fine-tuning can then be applied to further optimise a pretrained network for a specific task, such as constructing structure-property relationships [31]. One of the drawbacks with a standard CNN is the requirement for a predefined task and labelled data to train the network. On the other hand, an unsupervised method for training a CNN is the variational autoencoder (VAE) [32]. VAEs are composed of a pair of CNNs, one for encoding the input and another that aims to reconstruct the input from the encoded representation. Once trained, the encoder portion of the VAE has been shown to provide suitable microstructure parameterisation for structure-property predictions [33, 34]. The advantage here is that the fingerprinting process is abstracted away from the predictive model, providing an application agnostic parameterisation for microstructure that can be fed into arbitrary downstream tasks, which is the main focus of this paper.
@@
The trained VAE is then used to generate a dataset of microstructural fingerprints, and DAMASK [40] is used to perform CP simulations of uniaxial tension applied to the corresponding VEs. The resulting fingerprints and volume-averaged stress responses at each strain increment are then used to train and validate a fully connected neural network surrogate for the CP model. There are a wide range of options for the surrogate model, but a fully connected architecture was selected as there are no limitations on suitable input size (unlike Gaussian process regression). The time-dependence of the loading is implicit in the model, as this is constant for each simulation, so a fixed number of output nodes could be used. It is important to note here that this is a simple model and is presented more as a proof of concept to illustrate the effectiveness of VAE encoding for parameterising microstructure for a downstream task which possesses microstructural dependence. To extend this work to operate on an arbitrary load path, recurrent neural networks (RNNs) and/or transformers would need to be considered. Therefore, in this case, the full pipeline consists of VE generation, VAE fingerprinting, CP simulations of uniaxial tension and, finally, a fully connected network to predict stress response. However, more generally, the surrogate model could be swapped out for a different predictive model that deals with some other material properties of interest, but the same trained VAE could be used without any additional training of that portion of the pipeline. The latent space constructed by the trained VAE is explored with dimensionality reduction via PCA for a visual aid to observe microstructural feature clustering throughout the latent space, and spectral clustering is used to further quantify the qualitative results from the PCA. Finally, we show that fingerprinting microstructure VEs in this way can provide a suitable descriptor for parameterising microstructure as inputs for a microstructure-dependent CP surrogate model.
@@ 2.1 Datasets
Several datasets of 3D single phase VEs were constructed using Dream3D [41]. Each VE is converted into a PyTorch [42] tensor with shape $(3,64,64,64)$, i.e. a $64^{3}$ VE with 3 channels at each voxel to describe crystal orientation as an Euler angles in the Bunge convention. The Euler angles are normalised to a range of $[0,1]$ for training purposes, by dividing by $(2\pi,\pi,2\pi)$ across the three channels respectively. We also use the DAMASK Python library [43] to map the Euler angles into the crystallographic fundamental zone to remove crystal symmetries as a preprocessing step. This requires prior knowledge about the crystal structure. We consider cubic and hexagonal symmetries here and, as such, the trained models do not extend beyond that. Most of the results presented herein use a cubic crystal structure. The hexagonal close-packed (HCP) structure is introduced for the surrogate model, as it provides a more anisotropic stress response to train against. There are several options for taking symmetry into account, but it was deemed that addressing crystallographic symmetries during preprocessing provided the most stability during training. Alternatively, symmetries could be considered in the loss function during training. However, we found that this caused numerical instabilities, due to enforcing a discontinuous loss function, and it can be expensive to compute at each epoch, due to the requirement for computing all symmetries at each epoch. In theory, all symmetric variants of the training data could be computed prior to training, but it is not practical to store them in memory due to the size of the dataset.
2.1.1 VAE Training Dataset
@@ 2.2.1 Variational Autoencoders (VAEs)
$z_{i}=\mu_{i}+e^{\frac{1}{2}\ln(\sigma_{i})\mathcal{N}(0,1)},$ (1)
@@ 2.2.3 Loss Functions
$\mathcal{L}_{\text{FFT}}(x,\hat{x})=\frac{1}{m}\sum_{j=1}^{m}\left[(\text{Im}\{\mathcal{F}(x)_{j}\}-\text{Im}\{\mathcal{F}(\hat{x})_{j}\})^{2}+(\text{Re}\{\mathcal{F}(x)_{j}\}-\text{Re}\{\mathcal{F}(\hat{x})_{j}\})^{2}\right],$ (3)
@@ 2.3 Crystal Plasticity (CP) Model
The open source CP simulation software DAMASK [40] was used alongside MatFlow [47] for managing high-concurrency workflows to generate the training data for the surrogate model. The workflow generates a VE using Dream3D, as described in Section 2.1.1, post-processes the output into a format readable by DAMASK and performs a CP simulation with uniaxial loading along the $x$-axis. Each VE is deformed to a strain of 0.01 at a rate of $1\times 10^{-3}$ over 20 increments, to characterize the plastic yielding behaviour of the VE. The VE is also passed through the trained VAE to construct a microstructural fingerprint in parallel with performing the CP simulation. Finally, the workflow output is processed to yield a dataset containing each VE’s fingerprint, and the associated volume-averaged stress at each strain increment. The workflow was executed to run 50,000 simulations, 48,401 of which converged to a solution, with some runs not converging due to the severe plastic anisotropy of HCP magnesium. This was split into a training set of 40,000, with the rest reserved for model validation. Each simulation was run across 4 cores on Intel Sapphire Rapids CPUs, with a wall-clock time of $\sim 10$ minutes per simulation. The volume-averaged stress responses were normalised using min-max scaling, across the whole dataset, to a range of [0, 1], prior to splitting the dataset into training and validation partitions. Single crystal parameters for a magnesium phenomenological CP model [48] were used to generate this dataset, as it provides a relatively large spread in stress response compared to copper due to higher anisotropy. The anisotropy comes from magnesium having a HCP crystal structure, and so the VAE was retrained, in the same manner as previously described and with the relevant fundamental zone, before being used for fingerprint construction.
@@ 2.4 Surrogate Model
As discussed, the VAE fingerprinting process is intentionally distinct from the surrogate model, with the aim being to construct an application agnostic parameterisation for microstructure, rather than training a fully end-to-end model from microstructure to stress response. Here, we consider a simple surrogate model for the uniaxial CP model described in Section 2.3, as a proof of concept for including VAE fingerprints as a parameterisation of microstructure for yielding a microstructure-dependent surrogate model. Figure 2 shows a schematic of the general pipeline.
Figure 2: Schematic of the full pipeline of the VAE fingerprinting process with an example surrogate model. The region highlighted in blue is interchangeable with other microstructure-dependent surrogate model options, whilst retaining the trained VAE and fingerprinting processes.
A fully connected neural network was constructed with an input layer size that matches the dimensionality of the fingerprints being operated on. This is followed by a hidden layer of size 512, with Leaky ReLU activation, and a final 21-dimensional layer, with sigmoid activation, to output the normalised stress response for each discrete strain increment. The Adam optimiser is used here again, operating on the mean-squared error between the prediction and true stress from the DAMASK simulations. A learning rate of $1\times 10^{-5}$ was used with a weight decay of $1\times 10^{-5}$ and a batch size of 128 was used to introduce some stochasticity into the training process, rather than backpropagating after a forward pass with the entire training set (which would have fit into memory with the resources available). This stochasticity is favourable here, as the training dataset is relatively small.
@@ 3.1.2 Network Depth Selection
ResNet12 $0.070\pm 0.006$ $0.070\pm 0.006$
ResNet18 $0.029\pm 0.003$ $0.031\pm 0.004$
ResNet34 $0.075\pm 0.004$ $0.075\pm 0.004$
@@
Despite the drop in BCE loss from ResNet18 to ResNet34, the mean misorientation is considerably lower for ResNet18. This confirms that, although the spectral term in the combined loss function is proposed as a method for regularising the latent space [46], it does also act as a reconstruction loss as well. Further analysis of the latent space, to confirm that is has been suitably regularised, is presented in Section 3.3.
@@ 3.2 Reconstruction Accuracy
The large and elongated grains are reconstructed with an accuracy comparable to the training data, as shown in Table 5. However, reconstruction in the case of the small grains is clearly not representative of the input VE. This is, in part, due to a lack of representation of small grains in the training data and, also, the fact that there is much more information contained in the VEs with smaller grains, relative to the training set, so there is less potential compression available. If smaller grains are of interest, or there is a higher variance in grain size across the dataset, relative length scales would need to be reevaluated and the size of the latent space would need to be increased to accommodate this additional diversity in the training set. To verify this, an additional dataset of 110,000 volume elements was generated, with an even split between volume elements with a mean ESD of 2.73 and 4.01 (smaller than the initial training dataset, which contained a mean ESD of 5.78). 100,000 were used to train the VAE from scratch, with various latent space sizes and 10,000 were reserved for validation. The loss behaviour throughout training for each of these architectures is shown in Figure 7(a). In the case where $|z|=256$, the validation loss diverges away from the training loss, showing clear signs of overfitting. Increasing the size of the latent space starts to yield smooth behaviour at $|z|=2048$, with no sign of overfitting and an increase in accuracy as the size of the latent space is further increased to $|z|=4096$. Figure 7(b) shows some example reconstructions on unseen data, which demonstrates that small grains can indeed be resolved, provided that they are suitably represented in the training data, although the smoothing effect at grain boundaries is more pronounced than with the previous dataset, due to the increased number of grain boundaries in the VE.
Figure 7: Results from retraining the VAE with a dataset containing volume elements comprising small grains comparable to the edge case described in Section 2.1.2. (a) shows the loss behaviour throughout training for three latent space sizes and (b) shows slices from some example reconstructions.
Generalisation up to the large grains and elongated grains from the training data is feasible due to these morphologies being obtainable by placing grains from the training data with equal orientations adjacent to each other, i.e. they can be considered as a subset of the distribution that the training data represents, whereas the small grain case is extrapolation away from the training dataset, which is notoriously challenging to deal with in machine learning models. Suitable modification of the training data and increasing the size of the latent space can aid in encoding smaller grains, but the mean misorientation on unseen data is still higher than the cases that deal with larger grains due to the increased number of grain boundaries and the smoothing effect of the VAE on reconstructions causing blurred grain boundaries.
Morphology $|z|$ Mean misorientation
Large grains 256 $0.033\pm 0.010$
Small grains 256 $0.063\pm 0.001$
Small grains 512 $0.063\pm 0.001$
Small grains 1024 $0.065\pm 0.001$
Small grains 2048 $0.064\pm 0.001$
Small grains 256 (retrained) $0.061\pm 0.001$
Small grains 512 (retrained) $0.056\pm 0.0008$
Small grains 1024 (retrained) $0.073\pm 0.002$
Small grains 2048 (retrained) $0.048\pm 0.0008$
Small grains 4096 (retrained) $0.045\pm 0.0007$
Elongated grains 256 $0.036\pm 0.009$
Table 5: Mean misorientations between input VEs and their corresponding reconstructions for 100 samples from each of the morphological feature edge cases. The cases that state $|z|$ (retrained) refer to cases for which a bespoke dataset was created, with modified morphologies that are representative of the edge case, and used to train the VAE from scratch. This is to indicate that the edge cases that performed poorly (small grains) can be accounted for by modifying the training dataset.
@@
As before, the mean misorientation was calculated across 100 samples for each of these edge cases and the results are provided in Table 6. The reconstruction accuracy here is comparable to the training data for the brass and copper examples, with a slight drop in accuracy for the copper case.
@@
Brass $0.044\pm 0.003$
Copper $0.061\pm 0.003$
Cube $0.029\pm 0.003$
Goss $0.032\pm 0.004$
Table 6: Mean misorientations between input VEs and their corresponding reconstructions for 100 samples from each of the textured edge cases, with $|z|=256$.